OpenStudy (anonymous):

how do u evaluate lne^1/3

zepdrix (zepdrix):

\[\large \log(a^b)=b*\log(a)\] Gabbyyyy first of all, you need to remember one of your logarithm rules! This might help us, yes? :o

OpenStudy (anonymous):

yes is the answer 8?

zepdrix (zepdrix):